OpenStudy (toxicsugar22):

According to the following reaction, how many moles of water will be formed upon the complete reaction of 21.5 grams of hydrogen peroxide (H2O2)? 2 H2O2 (aq) = 2 H2O (l) + O2 (g)

OpenStudy (anonymous):

the same procedure just this time no need for limiting reagent based on the direction of the reaction just the same procedure lets start moles of H2O2=21.5/MM H2O2(34) mole ratio again 2H2O2=2H2O,THEN 21.5/(MMH2O2)=X MOLES H2O when found again moles =mass/molar mass moles of H20 ABOVE=Mass/MMH2O THEREFOR MASS=MM*MOLES MMH2O=18 NB;THIS IS BECAUSE WHEN THE MOLE RATIO HAS BEEN USED IN CALCULATING THE MOLES NO NEED TO MULTIPLY IT BY THE MOLAR MASS ANYMORE,I KEEP SAYING THIS BECAUSE IT USED TO BE CONFUSING FOR ME

OpenStudy (anonymous):

11.3823

OpenStudy (anonymous):

absolutely